ADNAN MUHAMMED

Grade 12,

A body is projected from the bottom of a smooth inclined plane with a velocity of 20 m/s. If it is just sufficient to carry it to the top in 4s, find the inclination and height of the plane. explain properly

A body is projected from the bottom of a smooth inclined
plane with a velocity of 20 m/s. If it is just sufficient to
carry it to the top in 4s, find the inclination and height of
the plane.   explain properly

Grade:10

1 Answers

Arun
4 years ago

let inclination be x

if time = 4 sec

20=mgsinx*4   (v=u+at)

mgsinx=5

l=80-8mgsinx    (s=ut+1\at2)

l=40

conserving energy

200=10lsinx

x=30 degree